# Curiosity-Driven Development of Action and Language in Robots Through Self-Exploration

### v4

@@ Abstract
Human infants acquire language and action gradually through development, achieving strong generalization from minimal experience, whereas large language models require exposure to billions of training tokens. What mechanisms underlie such efficient developmental learning in humans? This study investigates this question through robot simulation experiments in which agents learn to perform actions associated with imperative sentences (e.g., push red cube) via curiosity-driven self-exploration. Our approach integrates the active inference framework with reinforcement learning, enabling intrinsically motivated developmental learning. The simulations reveal several key findings: i) Generalization improves markedly as the scale of compositional elements increases. ii) Curiosity combined with motor noise yields substantially better learning than exploration without curiosity. iii) Rote pairing of sentences and actions precedes the emergence of compositional generalization. iv) Simpler, prerequisite-like actions develop earlier than more complex actions that depend on them. v) When exception-handling rules were introduced—where certain imperative sentences required executing inconsistent actions—the robots successfully acquired these exceptions through exploration and displayed a U-shaped performance curve characteristic of representational redescription in child language learning. Together, these results suggest that curiosity-driven exploration and active inference provide a powerful account of how intrinsic motivation and hierarchical sensorimotor learning can jointly support scalable compositional generalization and exception handling in both humans and artificial agents.
@@ Introduction
From the perspective of developmental psychology, infants acquire language through rich interaction with their embodied environments. Tomasello’s “verb-island” hypothesis argues that children initially learn verbs in specific, isolated contexts before generalizing across broader linguistic structures (?). He also emphasized the importance of embodiment in language acquisition, suggesting that grounding linguistic symbols in sensorimotor experiences is fundamental to language learning (?). This view aligns with other studies in developmental psychology highlighting the role of compositionality and generalization in language acquisition (?, ?, ?).
In linguistic terms, compositionality refers to the ability to construct novel configurations by systematically combining elements such as verbs, adjectives, and nouns. Generalization enables infants to apply learned components flexibly, allowing for the production and interpretation of utterances that have not been directly encountered previously. Although the number of possible compositions grows multiplicatively with the vocabulary size (i.e., number of verbs $\times$ number of adjectives $\times$ number of nouns), infants achieve generalization after experiencing only a small subset of learning examples. This suggests that the effective sample complexity could be proportional to the sum of elements rather than their product. This phenomenon is closely related to the “poverty of the stimulus” problem articulated by Chomsky (?), which asks how learners generalize so effectively given severely sparse input.
To investigate these mechanisms, one promising approach is to reconstruct developmental learning processes in machines and robots. The field of developmental robotics has long pursued this line of research, aiming to replicate human-like learning trajectories in embodied systems (?, ?, ?, ?). However, relatively few studies have focused on development of language and motor control under conditions of stimulus poverty. Existing work has primarily examined associative mappings between linguistic input and motor commands in one-shot or supervised batch learning schemes (?, ?, ?, ?). These approaches neglect the self-directed, developmental context of infant learning.
@@
A simulated mobile robot equipped with a manipulator arm, vision sensor, and distributed tactile sensors learns to generate motor movements in response to imperative sentences presented during each trial. These sentences are systematically composed of verbs, adjectives, and nouns, enabling evaluation of generalization performance under different levels of compositional complexity.
Beyond these primary developmental processes, we further explore the capacity for exception handling—a hallmark of flexible cognition. In human development, exceptions such as irregular verbs or inconsistent mappings often produce non-monotonic, U-shaped learning trajectories: children first apply a correct form, then overgeneralize it (producing errors), and finally recover the correct rule. This pattern has been widely interpreted as evidence of internal representational reorganization or representational redescription (?). Computationally, such U-shaped performance has been demonstrated in models of language acquisition and rule learning (?, ?, ?, ?, ?). Developmentally, these phenomena reflect the tension between rote memorization, generalization, and the later refinement of exception rules.
To examine whether similar mechanisms can emerge in artificial developmental systems, we hypothesize that robots trained through curiosity-driven active inference can also acquire exception-handling rules through exploratory learning. Specifically, we predict that when presented with a small number of abnormal linguistic rules—commands requiring reversed or inconsistent actions—the robot will exhibit a U-shaped performance curve: initial correct learning, followed by a decline due to overgeneralization, and eventual recovery as exception-specific representations are internalized. This pattern would indicate that the system develops not merely through associative reinforcement, but through representational reorganization analogous to cognitive processes observed in human infants.
@@
This study therefore tested the following hypotheses through simulation experiments: H1: Generalization performance improves as the scale of compositionality in the task increases. H2: Curiosity combined with motor entropy enhances the performance of developmental learning. H3: In the early phase, actions are generated only for exactly learned imperative sentences, but in later phases, the system generalizes to novel, unlearned compositions. H4: Primitive actions are acquired earlier, followed by more complex, prerequisite-dependent actions. H5: Exception-handling rules can be acquired through exploratory learning, exhibiting U-shaped developmental performance similar to that observed in human cognition.
@@ Results
The robot can maneuver by controlling velocity of left and right wheels independently, and also can move its arm by controlling rotation velocity of the yaw and pitch joint angles for acting on the objects. A camera with 16 x 16 pixels was fixed to the body for visual sensation. 16 touch sensors were distributed in the body and the arm, respectively, and rotation angles for the yaw and pitch were sensed as proprioception.
@@
At each step, the robot receives visual sensation, proprioception for the arm, tactile sensation , and two types of voices: the command voice and the tutor-feedback voice. The command voice takes the format of the imperative sentence described previously, and it comes every step continuously from the beginning. On the other hand, the feedback voice arrives whenever the robot achieves one of possible goals even if the achieved goal is not the imperative sentence told by the command voice, and it informs which goal has been achieved actually in the same format with the command voice. This potentially enhances the forward model to learn about own action as associated with linguistic representation. Finally, when the goal specified by the command voice is achieved, a reward is provided. Each trial episode ran for 30 steps, or terminated when the specified goal is achieved.
@@
The experiment examined effects of different levels of curiosity to the developmental learning processes using the basic setup. In the basic setup, full compositions of words (Table ) were used to generate the imperative sentences. However, the training was conducted using only 60 imperative sentences (33 percentage) out of 180 possible sentences. 120 untrained sentences were used for generalization test. For ten robots with different random seeds, the complete developmental learning process was iterated for 60000 epochs. The generalization test with unlearned imperative sentences was conducted for every 50 epochs.
The experiment was conducted by changing the levels of curiosity. Since the random latent variables are computed separately for each sensory modality, the complexity or curiosity can be computed for each sensory modality. Three levels of curiosity were considered in computing expected free energy $G$ wherein no curiosity: the curiosity terms for all sensory modalities are not included, sensory-motor curiosity: the curiosity terms only for vision, tactile sensation, and proprioception are included, all curiosity: the curiosity terms for all sensory modalities including feedback voice are included.
Fig. 2 shows the development of the generalization test performances in terms of success rate for goals specified by unlearned imperative sentences which are plotted for different action categories with different levels of curiosity. Shades areas represent 99% confidence intervals.
@@
The plots shows that the performance was improved significantly as the curiosity level was increased. Especially, the case of all actions with the all curiosity level shows the average success rate for unlearned goals reached a quite high value of 85 percentage even though the learning was conducted only for 33 percentage of all possible compositions.
It can be also seen in Fig. 2 that some action categories developed faster than other action categories. Especially in the all curiosity case, “watch” developed the fastest, “be near” did as the second, “touch the top” as the third, and “push left” “push right”, and “push forward” developed much later. This implies that simpler prerequisite-type actions develop earlier, and more complex actions requiring those prerequisite actions develop later. Actually, an action of watching an object should be prerequisite for all other object-targeted actions including an action of moving near by an object, which should be prerequisite again for actions of directly manipulating an object like pushing left/right an object or touching the top of it. Our observation accords with this.
@@
Video Experiment 2 shows examples of the behaviors of robots with the all curiosity level. It can be seen that in the intermediate phase of development, the robot often acts with play-like behavior without achieving specified goals while it quickly and accurately achieves its goals in the final phase of development.
@@
Some analysis were conducted for the purpose of gaining comprehension of the internal representation developed. We applied Principle Component Analysis (PCA) to the estimated posterior latent states corresponding to the command voice input, incorporating the verbs and adjectives stated in each command. Fig. 3 presents evidence that robots with all curiosity developed a compositional and generalizable understanding of these goals. At the midpoint of training, the latent representations begin to show consistent grouping of verbs and adjectives. For example, the verbs “watch,” “be near,” and “push forward” are tightly grouped, suggesting that these verbs are interpreted as similar. In contrast, the verb “push right” appears heavily separated from other verbs, and the verbs “touch the top” and “push left” also appear as distinct categories. After training, these clusters of verbs are more compact and separated. Within these clusters, there is loose sub-structuring by color: green and yellow tend to be on the left side of the cluster, while blue and magenta are on the right side of the cluster. Therefore, it can be said that each cluster represents a distinct linguistic concept while exhibiting relationship with others since the learning of visuo-proprioceptive-motor also contributes to this structuring. In contrast, Fig. 4 shows the same type of PCA results for a robot with no curiosity. “Be near,” “push forward,” and “push right” are heavily entangled with each other. Thus supports the idea that curiosity aids verb disentanglement and compositional learning.
@@
Fig. 4: PCA for language latent variables in the case of no curiosity. PCA applied to latent representations of command voice inputs after complete development.
@@
Next experiment examines the effects of scales of compositionality in learned examples to the generalization performance. For this purpose, experiments were conducted using reduced number of words for generating imperative sentences. While the previous basic setup used sentences composed of 6 verbs, 6 adjectives, and 5 object nouns as the full scale case, the middle scale case was prepared with 5 verbs, 5 adjectives, and 4 object nouns, and the small scale case with 4 verbs, 4 adjectives, and 3 object nouns. The exact words used for each setup are listed in Table 2. For all scaling cases, again only one third was used for learning examples while remained two third was used for generalization test. Other experimental conditions were also set as the same as the Experiment 1.
The experimental results are shown in Fig. 6. Shades areas represent 99% confidence intervals. It can be seen that although the learned goal test cases show equally high performance for all scales of compositionality, the generalization test for unlearned goal case shows that the success rate in the final trial decreases significantly (85 percentage to 25 percentage) as the compositionality scale decreases. This indicates that the generalization performance severely depends on the scale of compositionality in learning examples.
Fig. 6: Rolling success-rates for learned and unlearned goals with different compositionality scales. (A) Agents trained with all six verbs, all six adjectives, and all five nouns. (B) Agents trained with five verbs, five adjectives, and three nouns. (C) Agents trained with four verbs, four adjectives, and three nouns.
@@
This experiment examined how robots can acquire exception handling rules through developmental learning. While most command–action mappings were preserved, the commands “watch magenta pillar” and “be near green pole” were swapped: success required performing the other goal, not the one commanded. These mismatches required the robot to override its learned generalized knowledge.
@@
This simulation experiment was conducted using the same model parameters with the ones used in the previous experiments with 10 robots with 60,000 epochs of developmental trials. At the end of development the average success rate among 10 robots was 84 percentage for the learned goals, 76 percentage for unlearned goals, and 50 percentage for the exception handling cases. The average success rate for the exception handling cases is not so high most likely due to over-generalization.
@@
Panel A in Fig. 7 shows the rolling success-rate for achieving the exception goals for each of 10 individual robots while Panel B shows the success rate for the same goals but without applying the exception handling rules. The final success rate for the exception handling case is diverse ranging from 15 percentage to 90 percentage as can be seen in Panel A. It was also discovered that 7 out of 10 robots trained with these exceptions exhibit characteristic U-shaped curves: early success, followed by a drop, and eventual recovery with higher success rate than the earlier one. In contrast, monotonic increase of success rate can be seen in all 10 individuals in the case of learning without the exception handling rules. Statistical comparisons (detailed in the Supplementary Text) confirm that U-shaped patterns are significantly more prevalent in the exception condition than in the control ($p=.0025$).
@@
In Fig. 8, PCA illustrates how the internal representations of of goals in the seventh robot in Fig. 7 A evolve in a manner consistent with the U-shaped success rates. In panel A, early in training, goal embeddings are muddled without clear structure, reflecting a learning phase with minimal generalization. In panel B, midway in training, the exception command “watch magenta pillar” is embedded near other “watch” goals, while “be near green pole” clusters with other “be near” goals, despite these associations being incorrect. These observation indicates that overgeneralization has occurred. In panel C, late in training, “watch magenta pillar” is now embedded near “be near” goals, and “be near green pole” near “watch” goals, indicating that the robot has correctly handled these exceptions as swapped pairs. Here, it can be said that the representational redescription took place in the course of developmental learning of exception handling rules.
@@
Fig. 7: Comparison of the performance curve with and without applying the exception handling rules. (A) The development of success rate in achieving two goals which are swapped as exceptions. Red vertical lines depict the peaks and valleys of the learning, as defined in supplementary text. (B) The development of success rate in achieving the same two goals without applying the exception handling rules.
@@
Fig. 8: PCA for language latent variables for an individual developed with the exception handling rules. (A) Plot at 20000 epochs, (B) plot at 32500 epochs, and (C) plot at 57500 epochs wherein the circled ”N” and circled ”W” denote the sentences applied with the exception handling

### v5

@@ Abstract
Infants acquire language with generalization from minimal experience, whereas large language models require billions of training tokens. What underlies efficient development in humans? We investigated this problem through experiments wherein robotic agents learn to perform actions associated with imperative sentences (e.g., push red cube) via curiosity-driven self-exploration. Our approach integrates active inference with reinforcement learning, enabling intrinsically motivated developmental learning. The simulations reveal key findings corresponding to observations in developmental psychology. i) Generalization improves drastically as the scale of compositional elements increases. ii) Curiosity improves learning through self-exploration. iii) Rote pairing of sentences and actions precedes compositional generalization. iv) Simpler actions develop before complex actions depending on them. v) Exception-handling induces U-shaped developmental performance, a pattern like representational redescription in child language learning. These results suggest that curiosity-driven active inference accounts for how intrinsically motivated sensorimotor–linguistic learning supports scalable compositional generalization and exception handling in humans and artificial agents.
@@ Introduction
From the perspective of developmental psychology, infants acquire language through rich interaction with their embodied environments. Tomasello’s “verb-island” hypothesis argues that children initially learn verbs in specific, isolated contexts before generalizing across broader linguistic structures with compositionality (?). He also emphasized the importance of embodiment in language acquisition, suggesting that grounding linguistic symbols in sensorimotor experiences is fundamental to language learning (?). This view aligns with other studies in developmental psychology highlighting the role of compositionality and generalization in language acquisition (?, ?, ?).
In linguistic terms, compositionality refers to the ability to construct novel configurations by systematically combining elements such as verbs, adjectives, and nouns. Generalization enables infants to apply learned components flexibly, allowing for the production and interpretation of utterances that have not been directly encountered previously. Although the number of possible compositions grows multiplicatively with the vocabulary size (i.e., number of verbs $\times$ number of adjectives $\times$ number of nouns), children achieve generalization after experiencing only a small subset of learning examples. This suggests that the effective sample complexity could be proportional to the sum of elements rather than their product. This phenomenon is closely related to the “poverty of the stimulus” problem articulated by Chomsky (?), which asks how learners generalize so effectively given severely sparse input.
Beyond these, it is well known that children can develop the capacity for exception-handling, a hallmark of flexible cognition. In human development, exceptions such as irregular verbs or inconsistent mappings often produce non-monotonic, U-shaped learning trajectories: children first apply a correct form, then overgeneralize it (producing errors), and finally recover the correct rule. This pattern has been widely interpreted as evidence of internal representational reorganization or representational redescription (?). Computationally, such U-shaped performance has been demonstrated in models of language acquisition and rule learning (?, ?, ?, ?, ?). Developmentally, these phenomena reflect the tension between rote memorization, generalization, and the later refinement of exception rules.
How can humans develop capacity for compositionality with generalization even with exception-handling through learning of sparse input? To investigate this question, one promising approach is to reconstruct developmental learning processes in machines and robots. The field of developmental robotics has long pursued this line of research, aiming to replicate human-like learning trajectories in embodied systems (?, ?, ?, ?). However, relatively few studies have focused on development of language and motor control under conditions of stimulus poverty. Existing work has primarily examined associative mappings between linguistic input and motor commands in one-shot or supervised batch learning schemes (?, ?, ?, ?). These approaches neglect the self-directed, developmental context of infant learning.
@@
A simulated mobile robot equipped with a manipulator arm, vision sensor, and distributed tactile sensors learns to generate motor movements in response to imperative sentences presented during each trial. These sentences are systematically composed of verbs, adjectives, and nouns, enabling evaluation of generalization performance under different levels of compositional complexity. The model architecture employed in this study is based on our previous work (?) with key modifications to accommodate multi-modal sensorimotor integration. (See details in the Materials and Methods section.) Fig. 1 presents the model architecture, which is composed of three main components: a forward model, an actor part, and a critic part.
Fig. 1: The proposed model architecture. The model consists of a forward model, actor, and critic.
The forward model learns to predict the next sensation $o_{t+1}$ based on the current sensation $o_{t}$ and the executed motor command $a_{t}$. The sensation includes pixel-based vision, tactile sensation, arm joint proprioception, and voice for the sentences. To address the hidden state problem and probabilistic nature of the environment, the prediction is performed contextually and stochastically using the random latent variables $z_{t}$ and deterministic ones $h^{q}_{t}$. Here, $p(z_{t})$ is the prior probabilistic distribution of the random latent variables before observing the current sensation while $q(z_{t})$ is the posterior one after observation of the current sensation caused by the current motor command execution, as will be detailed later. The deterministic variables were shared for each sensory modality, while the random latent variables were allocated separately. This separation of random latent variables was necessary for the system to deal with different types of sensory modalities simultaneously. The actor module generates the next motor command $a_{t+1}$ based on the deterministic latent variable $h^{q}_{t}$, which integrates current sensation $o_{t}$. Based on $a_{t}$ and $h^{q}_{t}$, the critic generates $\widehat{Q}$, which is a prediction of the $Q$ value defined with Eq. The Employed Model. While the critic learns to make accurate predictions about future rewards in $Q$, the actor learns to produce motor commands which maximize the critic’s predictions $\widehat{Q}$.
The overall flow is: with an imperative sentence given by the tutor, the robot attempts to achieve the specified goal by generating a sequence of motor commands. Meanwhile, the forward model predicts the next sensation by inferring the posterior probability distribution $q(z_{t}|o_{t},h_{t-1})$ of the random latent variable $z$. This inference is conducted by minimizing the evidence free energy $F$ (Eq. 1). This consists of the complexity term represented by Kullback–Leibler divergence (KLD) between the estimated posterior and the prior, and the accuracy term as shown in the free energy principle (FEP). (See more details on the evidence free energy and expected free energy in the “Free Energy Principle, Active Inference, and Kawahara Model” section of the Supplementary Materials.)
$F_{\psi,t}=\underbrace{D_{KL}[q(z_{t}|o_{t},h_{t-1})||p(z_{t}|h_{t-1})]}_{\text{Complexity}}-\underbrace{\mathbb{E}_{q(z_{t})}[\log p(o_{t+1}|h_{t})]}_{\text{Accuracy}}.$ (1)
The forward model is trained iteratively by optimizing its learning parameters $\psi$ in the direction of minimizing the evidence free energy. The actor learns to generate motor command sequences in the direction of minimizing the expected free energy $G$ (Eq. 2) through reinforcement learning. This consists of the complexity term, extrinsic reward term, and the motor entropy term.
$G(a_{t})$ $=-\underbrace{D_{KL}[q(z_{t}|o_{t},h_{t-1})||p(z_{t}|h_{t-1})]}_{\text{complexity}}-\underbrace{r(s_{t},a_{t})}_{\text{Extrinsic Reward}}-\underbrace{\mathcal{H}(\pi_{\phi}(a_{t}|h_{t-1}))}_{\text{Entropy}}$ (2)
It is interesting to note that minimizing evidence free energy $F$ minimizes the complexity term, while minimizing expected free energy $G$ maximize the same complexity term. This means that motor commands are generated in the direction of maximizing the information gain attained after execution of the motor command, which is represented by KLD between the estimated posterior and the prior. This generates curiosity-driven exploration wherein the agent seeks out previously unencountered sensorimotor experiences. On the other hand, the learning parameters in the forward model is updated in the direction of minimizing the same complexity term, representing the latent conflict that is generated by novel experiences encountered during curiosity-driven exploration. Therefore, both processes of self-exploration and the forward model learning are racing each other, as if in competition.
@@
This study therefore tested the following hypotheses through simulation experiments: H1: Generalization performance improves drastically as the scale of compositionality in the task increases. H2: Curiosity combined with motor entropy enhances the performance of developmental learning. H3: In the early phase, actions are generated only for exactly learned imperative sentences, but in later phases, the system generalizes to novel, unlearned compositions. H4: Primitive actions are acquired earlier, followed by more complex, prerequisite-dependent actions. H5: Exception-handling rules can be acquired through exploratory learning, exhibiting U-shaped developmental performance similar to that observed in human development.
@@ Results
The robot can maneuver by controlling velocity of left and right wheels independently, and also can move its arm by controlling rotation velocity of the yaw and pitch joint angles for acting on the objects. A camera with 16 x 16 pixels was fixed to the body for visual sensation. 16 touch sensors were distributed in the body and the arm, and rotation angles for the yaw and pitch were sensed as proprioception.
@@
At each step, the robot receives visual sensation, proprioception for the arm, tactile sensation, and two types of voices: the command voice and the tutor-feedback voice. The command voice takes the format of the imperative sentence described previously, and it comes every step continuously from the beginning. On the other hand, the feedback voice arrives whenever the robot achieves one of possible goals even if the achieved goal is not the imperative sentence told by the command voice, and it informs which goal has been achieved actually in the same format with the command voice. This potentially enhances the forward model to learn about its own action as associated with linguistic representation. Finally, when the goal specified by the command voice is achieved, a reward is provided. Each trial episode ran for 30 steps, or is terminated when the specified goal is achieved.
@@
This experiment examined effects of different levels of curiosity to the developmental learning processes using the basic setup. In the basic setup, full compositions of words (Table ) were used to generate the imperative sentences. However, the training was conducted using only 60 imperative sentences (33 percentage) out of 180 possible sentences. 120 untrained sentences were used for generalization test. For ten robots with different random seeds, the complete developmental learning process was iterated for 60000 epochs. The generalization test with unlearned imperative sentences was conducted every 50 epochs.
The experiment was conducted by changing the levels of curiosity. Since the random latent variables are computed separately for each sensory modality, the complexity or curiosity can be computed for each sensory modality. Three levels of curiosity were considered in computing expected free energy $G$: no curiosity, wherein none of the curiosity terms for sensory modalities are included; sensory-motor curiosity, wherein the curiosity terms only for vision, tactile sensation, and proprioception are included; and all curiosity, wherein the curiosity terms for all sensory modalities including feedback voice are included.
Fig. 3 shows the development of the generalization test performances in terms of success rate for goals specified by unlearned imperative sentences, which are plotted for different action categories with different levels of curiosity. Shaded areas represent 99% confidence intervals. The plots show that the performance was improved significantly as the curiosity level was increased. Importantly, the case of all actions with the all curiosity level shows the average success rate for unlearned goals reached a quite high value of 85 percentage even though the learning was conducted only for 33 percentage of all possible compositions. It can be also seen in Fig. 3 that some action categories developed faster than other action categories. Especially in the all curiosity case, “watch” developed the fastest, “be near” did as the second, “touch the top” as the third, and “push left” “push right”, and “push forward” developed much later. This implies that simpler prerequisite-type actions develop earlier, and more complex actions requiring those prerequisite actions develop later. For example, an action of watching an object should be a prerequisite for all other object-targeted actions including an action of moving nearby an object, which should also be a prerequisite for actions of directly manipulating an object like pushing left/right an object or touching the top of it. Our observation accords with this.
@@
Video Experiment 2 shows examples of the behaviors of robots with the all curiosity level. It can be seen that in the intermediate phase of development, the robot often acts with play-like behavior without achieving specified goals. In the final phase of development, it quickly and accurately achieves its goals.
@@
Some analysis were conducted for the purpose of gaining comprehension of the internal representation developed. We applied Principle Component Analysis (PCA) to the estimated posterior latent states corresponding to the command voice input, incorporating the verbs and adjectives stated in each command. Fig. 4 shows an analysis of the internal representation which clarifies how robots with all curiosity developed a compositional and generalizable understanding of these goals. At the midpoint of training, the latent representations begin to show consistent grouping of verbs and adjectives. For example, the verbs “watch,” “be near,” and “push forward” are tightly grouped, suggesting that these verbs are interpreted as similar. In contrast, the verb “push right” appears heavily separated from other verbs, and the verbs “touch the top” and “push left” also appear as distinct categories. After training, these clusters of verbs are more compact and separated. Within these clusters, there is loose sub-structuring by color: green and yellow tend to be on the left side of the cluster, while blue and magenta are on the right side of the cluster. Therefore, it can be said that each cluster represents a distinct linguistic concept while exhibiting relationship with others since the learning of visuo-proprioceptive-motor also contributes to this structuring. In contrast, Fig. S5 shows the same type of PCA results for a robot with no curiosity. “Be near,” “push forward,” and “push right” are heavily entangled with each other. Thus supports the idea that curiosity aids verb disentanglement and compositional learning.
@@
This experiment examines the effects of scales of compositionality in learned examples to the generalization performance. For this purpose, experiments were conducted using a reduced number of words for generating imperative sentences. While the previous basic setup used sentences composed of 6 verbs, 6 adjectives, and 5 object nouns as the full scale case, the middle scale case was prepared with 5 verbs, 5 adjectives, and 4 object nouns, and the small scale case with 4 verbs, 4 adjectives, and 3 object nouns. The exact words used for each setup are listed in Table 2. For all scaling cases, only one third was used for learning examples while the remaining two thirds were used for generalization test. Other experimental conditions were the same as in Experiment 1.
The experimental results are shown in Fig. 6. Shaded areas represent 99% confidence intervals. It can be seen that although the learned goal test cases show equally high performance for all scales of compositionality, the generalization test for unlearned goal case shows that the success rate in the final trial decreases significantly (85 percentage to 25 percentage) as the compositionality scale decreases. This indicates that the generalization performance severely depends on the scale of compositionality in learning examples.
Fig. 6: Rolling success rates for learned and unlearned goals with different compositionality scales. (A) Agents trained with all six verbs, all six adjectives, and all five nouns. (B) Agents trained with five verbs, five adjectives, and three nouns. (C) Agents trained with four verbs, four adjectives, and three nouns.
@@
This experiment examined how robots can acquire exception-handling rules through developmental learning. While most command–action mappings were preserved, the commands “watch magenta pillar” and “be near green pole” were swapped: success required performing the other goal, not the one commanded. These mismatches required the robot to override its learned generalized knowledge.
@@
This simulation experiment was conducted using the same model parameters used in the previous experiments with 10 robots with 60,000 epochs of developmental trials. At the end of development the average success rate among 10 robots was 85 percentage for the learned goals, 75 percentage for unlearned goals, and 49 percentage for the exception-handling cases. The average success rate for the exception-handling cases is not so high (only some individuals are successful) which should be reasonable by considering the task complexity.
@@
Panel A in Fig. 7 shows the rolling success rate for achieving the exception goals for each of 10 individual robots while Panel B shows the success rate for the same goals but without applying the exception-handling rules. The final success rate for the exception-handling case is diverse, ranging from 15 percentage to 90 percentage as can be seen in Panel A. It was also discovered that 7 out of 10 robots trained with these exceptions exhibit characteristic U-shaped curves: early success, followed by a drop, and eventual recovery with higher success rate than the earlier one. In contrast, monotonic increase of success rate can be seen in all 10 individuals in the case of learning without the exception-handling rules. Statistical comparisons (detailed in the Supplementary Text) confirm that U-shaped patterns are significantly more prevalent in the exception condition than in the control ($p=.0025$).
@@
In Fig. 8, PCA illustrates how the internal representations of goals in the seventh robot in Fig. 7 (A) evolve in a manner consistent with the U-shaped success rates. In panel A, early in training, goal embeddings are muddled without clear structure, reflecting a learning phase with minimal generalization. In panel B, midway in training, the exception command “watch magenta pillar” is embedded with other “watch” goals, while “be near green pole” clusters is embedded with other “be near” goals, despite these associations being incorrect. These observations indicate that overgeneralization has occurred. In panel C, late in training, “watch magenta pillar” is now embedded near “be near” goals, and “be near green pole” near “watch” goals, indicating that the robot has correctly handled these exceptions as swapped pairs. Here, it can be said that the representational redescription took place in the course of developmental learning of exception-handling rules.
@@
Fig. 7: Comparison of the performance curve with and without applying the exception handling rules for 10 individuals. (A) The development of success rate in achieving two goals which are swapped as exceptions. Red vertical lines depict the peaks and valleys of the learning, as defined in supplementary text. (B) The development of success rate in achieving the same two goals without applying the exception handling rules.
@@
Fig. 8: PCA for language latent variables for an individual developed with the exception handling rules. (A) Plot at 20000 epochs, (B) plot at 32500 epochs, and (C) plot at 57500 epochs wherein the circled “N” and circled “W” denote the sentences applied with the exception handling.